# Can LLMs Maintain Fundamental Abilities under KV Cache Compression?

### v1

@@ Abstract
This paper investigates an under-explored challenge in large language models (LLMs): the impact of KV cache compression methods on LLMs’ fundamental capabilities. While existing methods achieve impressive compression ratios on long-context benchmarks, their effects on core model capabilities remain understudied. We present a comprehensive empirical study evaluating prominent KV cache compression methods across diverse tasks, spanning world knowledge, commonsense reasoning, arithmetic reasoning, code generation, safety, and long-context understanding and generation.Our analysis reveals that KV cache compression methods exhibit task-specific performance degradation. Arithmetic reasoning tasks prove particularly sensitive to aggressive compression, with different methods showing performance drops of $17.4\%$-$43.3\%$. Notably, the DeepSeek R1 Distill model exhibits more robust compression tolerance compared to instruction-tuned models, showing only $9.67\%$-$25.53\%$ performance degradation. Based on our analysis of attention patterns and cross-task compression performance, we propose ShotKV, a novel compression approach that distinctly handles prefill and decoding phases while maintaining shot-level semantic coherence. Empirical results show that ShotKV achieves $9\%$-$18\%$ performance improvements on long-context generation tasks under aggressive compression ratios.
[…]
@@ 1 Introduction
The emergence of Key-Value (KV) cache compression techniques has become crucial for efficient LLM deployment, primarily due to the increasing demands of memory management during inference. This need arises from the significant evolution of Large Language Models (LLMs), which now excel at processing extensive documents spanning thousands of tokens for tasks like question answering and summarization (Raffel et al., 2020; Brown et al., 2020; Chowdhery et al., 2022; Tay et al., 2022; Touvron et al., 2023a, b). The ability to handle such lengthy inputs has been transformed by two concurrent developments: breakthroughs in ML system architectures for processing long sequences (Dao et al., 2022; Dao, 2024; Jacobs et al., 2023; Xiao et al., 2024), and innovations in model design (Chen et al., 2023a; Xiong et al., 2024; Chen et al., 2023b; Peng et al., 2024). However, this enhanced capability comes with significant computational costs; as context lengths grow, the GPU memory requirements for inference operations increase substantially (AI21, 2024; X.AI, 2024; Reid et al., 2024; Anthropic, 2024; DeepSeek-AI, 2024; Liu et al., 2024a). This challenge has made the development of efficient KV cache compression strategies a critical focus in the field of LLM deployment and optimization.
Multiple studies have proposed innovative approaches to address this challenge through selective token retention strategies (Xiao et al., 2024; Zhang et al., 2023; Li et al., 2024b; Ge et al., 2023; Cai et al., 2024; Fu et al., 2024; Yang et al., 2024; Adnan et al., 2024; Liu et al., 2024e; Tang et al., 2024; Liu et al., 2025). Pioneering works such as H2O (Zhang et al., 2023) and SnapKV (Li et al., 2024b) demonstrate that retaining approximately 50% of KV cache entries can maintain model performance while achieving substantial memory efficiency. However, current research primarily evaluates these methods on long-context scenarios like LongBench (Bai et al., 2023, 2025) and Need-In-A-Haystack (NIAH) (Kamradt, 2023), leaving their broader impact on fundamental LLM capabilities largely unexplored. Our preliminary analysis, as shown in Figure 1, reveals two critical findings: (a) arithmetic reasoning tasks suffer significantly higher performance degradation under compression compared to long-context tasks, and (b) attention patterns in long-context scenarios exhibit notably higher sparsity than arithmetic tasks. These observations suggest that current evaluation frameworks, which focus predominantly on long-context performance, may inadequately capture the full spectrum of impacts on model capabilities. Further analysis of attention patterns, visualized in Figure 2, reveals distinct task-specific behaviors: while world knowledge and commonsense reasoning tasks exhibit universal attention distributions, arithmetic reasoning and safety tasks demonstrate more specialized patterns. Specifically, arithmetic reasoning tasks display increased attention sparsity, indicating focused attention on individual prompt examples, while safety tasks show concentrated attention on the system prompt. In contrast, world knowledge and commonsense reasoning tasks demonstrate more uniform attention distribution across the entire prompt. These varying attention patterns—visualized through colored squares representing system prompts, shot examples, and questions—provide insights into task-specific context utilization and motivate our investigation into how compression affects various factors including model size, prompt length, and task type.
Figure 2: Attention heatmap on different tasks. The heatmap is generated by the attention scores of the 15-th layer of the LLaMA-3.1-8B-Instruct attention head 10.
Motivated by these initial findings, we conduct a systematic study to comprehensively evaluate KV cache compression methods across different tasks (detailed in Section 4). Our analysis reveals several key findings: (1) Performance degradation is highly task-dependent, with arithmetic reasoning tasks showing particular sensitivity to aggressive compression; (2) Multi-step reasoning LLMs (O1 and R1) demonstrate higher compression robustness compared to instruction-tuned models; (3) Shorter prompts are more vulnerable to compression effects; (4) Chunk-level compression strategies show superior performance on complex long-context reasoning tasks; (5) Tasks with larger prompt-based performance gains exhibit higher compression sensitivity; and (6) Long-context generation tasks are particularly sensitive to compression. These findings provide valuable insights into the relationship between compression methods and model capabilities, motivating our development of ShotKV.
@@
We hope our work can provide the research community with insightful perspectives on KV cache compression’s impact on LLMs. To the best of our knowledge, we are the first to comprehensively evaluate and analyze the impact of KV cache compression on LLMs’ fundamental abilities. Our main contributions are summarized as follows:
  • Through comprehensive evaluation of KV cache compression methods across diverse tasks and compression ratios, we demonstrate that task-specific sensitivity to compression varies significantly, with performance degradation ranging from 1% to 40% under aggressive compression settings and multi-step reasoning LLM (DeepSeek R1) is more robust than instruction-tuned LLM.
@@
  • Our systematic investigation reveals multiple critical factors influencing compression sensitivity, including model training dynamics, prompt length characteristics, task-specific requirements, long-context reasoning and long-context generation capabilities. These findings provide generalizable insights applicable across various compression methodologies and model architectures.
@@
  • We introduce ShotKV, an innovative compression framework that distinctively manages prefill and decoding phases while maintaining shot-level semantic integrity. This approach demonstrates marked improvements in performance, particularly on complex long-context arithmetic reasoning and long-context generation tasks.
2 Related Works
2.1 Key–value Cache Optimization Techniques
KV cache is the core component in LLM inference, which avoids repetitive computations by caching Key and Value vectors. However, the cost of caching KV increases exponentially with the expansions of the model size and context length (Pope et al., 2023). Some approaches have been published to alleviate the issue. For instance, KV Compression designs efficient content selection strategies to filter and manage tokens (Zhang et al., 2023; Adnan et al., 2024). Some methods identify important tokens by focusing on high attention allocation (Li et al., 2024b), while others optimize token selection by combining attention scores with value vector norms to improve importance evaluation (Guo et al., 2024). Techniques like PyramidInfer reduce critical contexts layer-by-layer based on the distribution of attention scores (Yang et al., 2024), and StreamingLLM preserves attention sinks to maintain stable performance in extended sequences (Xiao et al., 2024). Researchers reduce storage costs by merging similar context representations and solving input disturbances caused by compression (Zhang et al., ). For instance, CaM (Zhang et al., ) works by integrating the KV cache to be dropped into the retained cache in proportion to the attention weight. In addition, Yao et al. (2024) proposes CacheBlend to achieve selective KV recompute. Only partial KVs of crucial tokens are updated to reduce the delay in the prefill stage and increase the throughput. Besides, the dynamic budget allocation method is also used to optimize the KV Cache, which adjusts the resource allocation in real-time according to the importance of the context, providing a balance between performance and efficiency in multi-task inference scenarios (Cai et al., 2024; Feng et al., 2024; Liu et al., 2025).
2.2 Evaluation of LLMs’ Fundamental Abilities
Accurately evaluating the fundamental capabilities of large language models is crucial for understanding their true potential and limitations. The evaluation typically spans across several key dimensions: world knowledge tasks like MMLU (Hendrycks et al., 2020),BBH (Suzgun et al., 2022) assess models’ grasp of diverse domains through multiple-choice questions; commonsense reasoning tasks such as CSQA (Talmor et al., 2019) evaluate inference and context understanding abilities; arithmetic reasoning benchmarks like GSM8K (Cobbe et al., 2021) test mathematical problem-solving capabilities through step-by-step reasoning; code generation tasks including HumanEval (Chen et al., 2021) measure the ability to generate functionally correct code; and safety evaluations using benchmarks like JailBreakV (Luo et al., 2024) assess models’ robustness against harmful content generation. Additionally, long-context benchmarks such as LongBench (Bai et al., 2023, 2025) and Need-In-A-Haystack (NIAH) (Kamradt, 2023) aiming to evaluate models’ long-context summarization and retrieval capabilities. Furthermore, LongGenBench (Liu et al., 2024d) evaluate models’ abilities to process and generate responses for extended input sequences. And recently, many-shot in-context learning has been recognized as a long-context reasoning paradigm (Agarwal et al., 2024), which considers the number of shots as a critical factor in the performance of LLMs. While these tasks typically employ automatic evaluation metrics for standardization, KV cache compression may introduce unique challenges, particularly in tasks requiring complex reasoning chains or extensive knowledge retrieval.
@@ 3 Preliminary
4 Experiments Design
4.1 Experimental Setups
In this section, we will introduce the experimental setups, including the datasets, models, and evaluation environment.
@@ 4.1 Experimental Setups
To evaluate the performance of KV cache compression on LLMs’ overarching capabilities, we assess five benchmark categories: World Knowledge using MMLU (Hendrycks et al., 2020), measured by accuracy; CommonSense Reasoning using CommonsenseQA (Talmor et al., 2019) , evaluated through multiple-choice accuracy; Arithmetic Reasoning using GSM8K (Cobbe et al., 2021), assessed by solve rate; Code Generation using HumanEval (Chen et al., 2021), measured by pass@1 rate on test cases; and Safety using JailBreakV (Luo et al., 2024), evaluated by attack success rate. Furthermore, we test the performance of KV cache compression on LongGenBench (Liu et al., 2024d), a long-context generation benchmark.
@@
We conduct experiments on a series of LLMs, including LLaMA-3.1-8B, LLaMA-3.1-8B-Instruct (Dubey et al., 2024), and multi-step reasoning LLM DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025).
@@
To comprehensively investigate the potential impact on KV cache compression methods, we select the following methods: StreamingLLM (Xiao et al., 2024), SnapKV (Li et al., 2024b), H2O (Zhang et al., 2023), PyramidKV (Cai et al., 2024), PyramidInfer (Yang et al., 2024), and ChunkKV (Liu et al., 2025).
@@
Evaluation Environment
We use the lm-evaluation-harness (Gao et al., 2023) library to load the models and evaluate the performance. The evaluation environment is a NVIDIA A40 GPU server.
4.2 Results and Analysis
In this section, we will present the results and analysis of the experiments. For comprehensive results, please refer to Section B.2.
@@ 4.2 Results and Analysis
Figure 4: Performance Comparison of KV Cache Compression Methods Across Tasks. Note: The y-axis scales vary across different tasks. Results for R1-GSM8K (e) were obtained using the DeepSeek-R1-Distill-Llama-8B model.
@@
KV cache compression methods show task-dependent performance degradation, with varying sensitivity thresholds across different benchmark categories. As demonstrated in Figure 3, all tasks maintain stable performance at compression ratios above 40%, but exhibit distinct degradation patterns below this threshold. GSM8K, HumanEval, and JailBreakV tasks demonstrate the highest compression sensitivity, characterized by precipitous performance declines. Figure 4 illustrates the detailed performance impact of various KV cache compression methods across different tasks. This degradation is most pronounced in GSM8K (d), where performance deteriorates significantly below 20% compression ratio, with accuracy dropping from approximately 0.75 to below 0.5. Among the evaluated methods, ChunkKV (Liu et al., 2025) and PyramidKV (Cai et al., 2024) consistently demonstrate superior stability across most tasks, while StreamingLLM (Xiao et al., 2024) exhibits heightened sensitivity to aggressive compression. Additionally, R1-GSM8K (e) indicates that R1 LLMs demonstrate enhanced robustness to KV cache compression.
Figure 5: Performance Comparison of KV Cache Compression Methods on different training dynamics on GSM8K
@@
Multi-step reasoning LLMs are more robust to KV cache compression. Figure 5 presents a comparative analysis of LLaMA-3.1-8B across its base (w/o instruct tuned), instruct-tuned, and DeepSeek-R1 distilled variants, illustrating their averaged performance across five compression methods with confidence intervals. Although all three variants exhibit performance degradation at low compression ratios, their degradation trajectories differ significantly. The R1 distilled model demonstrates superior stability, maintaining performance around 0.60 even at a 10% compression ratio. While the instruct-tuned model achieves a higher initial accuracy (0.8), it exhibits heightened compression sensitivity, with performance deterioration initiating at 30% compression ratio and declining sharply to approximately 0.5 at 10% ratio. These findings suggest that while multi-step reasoning LLMs demonstrate enhanced robustness to KV cache compression, and instruct-tuning improves overall model performance, the latter may inadvertently increase model vulnerability to aggressive compression, particularly at compression ratios below 30%.
@@
Observation 3.
Short prompt length is more sensitive to KV cache compression. As shown in Figure 6, the impact of KV cache compression varies significantly with different prompt lengths (shot numbers). One-shot and two-shot scenarios demonstrate higher sensitivity to compression, with performance dropping more sharply below 30% compression ratio compared to scenarios with more shots (4-8). For instance, in 1-shot settings, performance decreases from 0.5 to 0.05 when compression ratio drops from 30% to 10%, while 8-shot settings show only drop from 0.75 to 0.5 reduction under the same conditions. This suggests that longer prompts provide more information that helps maintain model performance even under aggressive compression. The increased robustness with higher shot numbers may be attributed to the model’s ability to leverage multiple examples, making it less dependent on perfect preservation of each individual example in the compressed KV cache.
Observation 4.
Chunk-level compression is more effective for long-context arithmetic reasoning tasks. Inspired by Agarwal et al. (2024), we consider many-shot in-context learning as a long-context reasoning task, which is more complex than existing long-context benchmarks, such as LongBench and NIAH. Figure 7 shows the performance of KV cache compression methods on a 50-shot GSM8K task, where the prompt length exceeds 4K tokens. From the figure, we observe that ChunkKV (Liu et al., 2025) demonstrates the most stability when the compression ratio is below 10% on both LLaMA-3.1-8B-Instruct and DeepSeek-R1-Distill-Llama-8B, indicating that in more complex long-context arithmetic reasoning tasks, chunk-level retention is more effective at preserving semantic information.
(a) Many-shot GSM8K on LLaMA3.1-8B-Instruct
(b) Many-shot GSM8K on DeepSeek-R1-Distill-Llama-8B
Figure 7: Many-shot scenario on KV cache compression
@@
Tasks with larger prompt-based performance gains show higher sensitivity to KV cache compression. As shown in Table 1, different tasks exhibit varying levels of performance improvement from zero-shot to few-shot prompting. GSM8K shows a dramatic improvement of 50.41%, while MMLU demonstrates a more modest gain of 6.20%. From Figure 3, we find that tasks with larger prompt-based improvements, such as GSM8K, are more sensitive to KV cache compression. This suggests that when a task heavily relies on in-context examples to achieve better performance, the compression of these crucial prompt elements has a more substantial impact on model performance. In contrast, tasks like MMLU, where the performance gain from prompting is smaller, show more resilience to KV cache compression, likely because the model relies more on its inherent knowledge rather than the specific examples in the prompt.
@@
GSM8K $0.2904$ $0.7945$ $+0.5041$
MMLU $0.6262$ $0.6882$ $+0.0620$
Observation 6.
KV cache compression exhibits significant performance degradation on Long-Context Generation tasks. As demonstrated in Table 2, our evaluation of three unified compression methods—StreamingLLM, H2O, and PyramidInfer—on LongGenBench-GSM8K reveals substantial performance limitations. On this arithmetic reasoning task with approximately 4k token generation length, the compression methods show notable deterioration, with performance declining by more than 20% at compression ratios below 30%. The ShotKV is our proposed method aiming to improve the performance of KV cache compression on Long-Context Generation tasks, details in Section 5.
Table 2: KV cache compression methods’ performance on LongGenBench-GSM8K

### v2

@@ Abstract
This paper investigates an underexplored challenge in large language models (LLMs): the impact of KV cache compression methods on LLMs’ fundamental capabilities. Although existing methods achieve impressive compression ratios on long-context benchmarks, their effects on core model capabilities remain understudied. We present a comprehensive benchmark KVFundaBench to systematically evaluate the effects of KV cache compression across diverse fundamental LLM capabilities, spanning world knowledge, commonsense reasoning, arithmetic reasoning, code generation, safety, and long-context understanding and generation.Our analysis reveals serval key findings: (1) Task-Dependent Degradation; (2) Model-Type Robustness (3) Prompt Length Vulnerability; (4) Chunk-Level Superiority; (5) Prompt-Gain Sensitivity; (6) Long-Context Generation Sensitivity. Based on our analysis of attention patterns and cross-task compression performance, we propose ShotKV, a novel compression approach that distinctly handles prefill and decoding phases while maintaining shot-level semantic coherence. Empirical results show that ShotKV achieves $9\%$-$18\%$ performance improvements on long-context generation tasks under aggressive compression ratios.
$\dagger$\$$\dagger$\$footnotetext: Corresponding Author.
@@ 1 Introduction
The evolution of Large Language Models (LLMs) to process large documents for tasks such as answering and summarizing questions [1, 2, 3, 4, 5, 6], spurred by breakthroughs in system architectures [7, 8, 9, 10] and model design [11, 12, 13, 14], has significantly increased GPU memory demands during inference [15, 16, 17, 18, 19, 20], making the development of efficient key value (KV) cache compression strategies a critical focus for LLM deployment and optimization.
To address this, numerous studies have proposed selective token retention strategies [10, 21, 22, 23, 24, 25, 26, 27, 28, 29], with pioneering works such as H2O [21] and SnapKV [22] showing that retaining approximately 50% of KV cache entries can balance model performance with significant memory savings. However, current research primarily evaluates these methods in long-context scenarios such as LongBench [30, 31] and Need-In-A-Haystack (NIAH) [32], leaving their impact on a spectrum of fundamental LLM capabilities—such as arithmetic reasoning, world knowledge, commonsense reasoning, and safety, particularly concerning their distinct attention patterns under compression, largely unexplored. To this end, we introduce KVFundaBench, a benchmark designed to systematically assess the effects of KV cache compression across these diverse fundamental capabilities and their underlying attention dynamics. The benchmark includes 5 categories of tasks: world knowledge, commonsense reasoning, arithmetic reasoning, code generation, and safety. Our comprehensive evaluations using KVFundaBench reveal several critical findings: we observe, as shown in Figure 1, that arithmetic reasoning tasks suffer significantly higher performance degradation under compression compared to long-context tasks, and that attention patterns in long-context scenarios exhibit notably higher sparsity. These initial results suggest that existing evaluation frameworks, which focus predominantly on long-context performance, may not adequately capture the full impact spectrum. Our KVFundaBench reveals several key findings: (1) Task-Dependent Degradation: Performance degradation is highly task-dependent, with arithmetic reasoning tasks showing particular sensitivity to aggressive compression; (2) Model-Type Robustness: Multi-step reasoning LLMs demonstrate higher compression robustness compared to instruction-tuned models; (3) Prompt Length Vulnerability: Shorter prompts are more vulnerable to compression effects; (4) Chunk-Level Superiority: Chunk-level compression strategies show superior performance on complex long-context reasoning tasks; (5) Prompt-Gain Sensitivity: Tasks with larger prompt-based performance gains exhibit higher compression sensitivity; and (6) Long-Context Generation Sensitivity: Long-context generation tasks are particularly sensitive to compression. These findings provide valuable insights into the relationship between compression methods and model capabilities, motivating our development of ShotKV, which is a new KV cache compression method with separate compression methods for prefill and decoding phases.
@@
We hope our work can provide the research community with insightful perspectives on the impact of KV cache compression on LLMs. Our main contributions are summarized as follows:
  • Introduce KVFundaBench to systematically evaluate the effects of KV cache compression across diverse fundamental LLM capabilities, we demonstrate that task-specific sensitivity to compression varies significantly, with performance degradation ranging from 1% to 40%.
@@
  • Our systematic investigation reveals multiple critical factors influencing compression sensitivity, including model training dynamics, prompt length characteristics, task-specific requirements, long-context reasoning, and long-context generation capabilities.
@@
  • We introduce ShotKV, an innovative compression framework that distinctively manages the prefill and decoding phases while maintaining the semantic integrity of the shot level.
@@ 2 Preliminary
3 Benchmark Design
3.1 Benchmark Setups
In this section, we will introduce the KVFundaBench setups, including the datasets, models, and evaluation environment.
Table 1: A summarized view of KVFundaBench
Dataset Task Type # Test
MMLU [33] World Knowledge 14,079
GSM8K [34] Arithmetic 1,319
CSQA [35] Commonsense 1,221
HumanEval [36] Code Generation 164
JailBreakV [37] Safety 28,000
LongGenBench [38] Long-Context Generation 23,000
@@ 3.1 Benchmark Setups
To evaluate the performance of KV cache compression on LLMs’ overarching capabilities, we assess five benchmark categories: World Knowledge (WK) using MMLU [33], measured by accuracy; CommonSense Reasoning (CSR) using CommonsenseQA [35] , evaluated through multiple-choice accuracy; Arithmetic Reasoning (AR) using GSM8K [34], assessed by solve rate; Code Generation (CG) using HumanEval [36], measured by pass@1 rate on test cases; and Safety (SA) using JailBreakV [37], evaluated by attack success rate. Furthermore, we test the performance of KV cache compression on LongGenBench [38], a long-context generation (LG) benchmark. Table 1 summarizes the statistics of the data set.
Figure 2: Attention heatmap on different tasks.
@@
We conduct experiments on a series of LLMs, including LLaMA-3.1-8B, LLaMA-3.1-8B-Instruct [39], Mistral-7B-Instruct [40], and multi-step reasoning LLM DeepSeek-R1-Distill-Llama-8B [41].
@@
To thoroughly investigate the potential impact on KV cache compression methods, we select the following methods: StreamingLLM [10], SnapKV [22], H2O [21], PyramidKV [24], PyramidInfer [26], and ChunkKV [42].
@@
3.2 Attention Pattern Analysis on KVFundaBench
(a) Attention Distribution with Sink Tokens
(b) Attention Distribution without Sink Tokens
Figure 3: Cumulative attention score distribution for Long-Context and Arithmetic. (a) Overall distribution including initial sink tokens, showing high initial concentration. (b) Distribution without sink tokens (first 4 tokens removed), revealing that Arithmetic’s non-sink attention is more diffuse compared to Long-Context’s.
To better understand the task-specific sensitivity to KV cache compression observed in our experiments, we analyze the attention patterns of LLMs on KVFundaBench. This analysis provides key insights into why different tasks respond differently to compression, from two perspectives: (1) task-specific attention patterns; and (2) attention distributions.
@@ 3.2 Attention Pattern Analysis on KVFundaBench
As shown in Figure 2, our benchmark reveals different attention behaviors in tasks: world knowledge and common sense reasoning tasks exhibit universal attention distributions, while arithmetic reasoning and safety tasks display more specialized patterns. In particular, arithmetic reasoning tasks show increased attention sparsity (i.e. focused attention on individual prompt examples), and safety tasks concentrate attention on system prompts, in contrast to the more uniform attention distribution seen in world knowledge and commonsense reasoning.
To further investigate the attention dynamics that might explain the task-specific sensitivities to KV cache compression, we analyzed the cumulative attention score distributions, as illustrated in Figure 3. Figure 3(a) depicts the overall attention distribution, which includes the initial sink tokens [10]. In this view, both long-context and arithmetic tasks demonstrate a very similar pattern: a steep initial rise where the top 1% of tokens capture over 60% of the total attention mass. This highlights the predominant role of sink tokens in attracting attention, regardless of the specific task.
However, a more distinct pattern emerges when these initial sink tokens (specifically, the first four tokens) are excluded from the analysis, as shown in Figure 3(b). Within the remaining non-sink tokens, the attention distribution for arithmetic tasks becomes notably more diffuse, with a slower accumulation of attention mass. For instance, the top 20% of non-sink tokens in arithmetic cover only about 37% of the attention within their own non-sink group. In contrast, long-context’s non-sink tokens exhibit a relatively more concentrated attention profile, where the top 20% of its non-sink tokens capture approximately 61.5% of the attention within their non-sink set. This divergence suggests that while sink tokens provide a common, strong attentional anchor, the subsequent distribution of attention across task-relevant (non-sink) tokens varies. The more diffuse attention in arithmetic’s non-sink tokens implies a reliance on a broader set of contextual cues for its structured reasoning, potentially making it more vulnerable when compression begins to impact these non-sink tokens.
These detailed analyses of attention distributions ( Figure 2 and Figure 3) reveal that LLMs engage different contextual information and attention strategies when performing long-context tasks versus tasks requiring fundamental abilities such as arithmetic reasoning. This highlights the necessity of evaluating KV cache compression beyond long-range dependencies to specifically assess its impact on diverse fundamental capabilities, owing to their distinct attentional mechanisms.
3.3 Results and Analysis
In this section, we present the results and an analysis of the experiments. For detailed results, see Section B.2.
Figure 5: Performance Comparison of KV Cache Compression Methods on KVFundaBench. Results for R1-AR (f) were obtained using the DeepSeek-R1-Distill-Llama-8B model. ShotKV is our proposed method; details can be found in Section 4.
Evaluation Environment
We use the lm-evaluation-harness [43] library to load the models and evaluate the performance. The evaluation environment is a NVIDIA A40 GPU server.
@@ 3.3 Results and Analysis
Task-Dependent Degradation: KV cache compression methods show task-dependent performance degradation, WK and CSR are more robust to KV cache compression. As demonstrated in Figure 4, all tasks maintain stable performance at compression ratios above 40%, but exhibit distinct degradation patterns below this threshold. Arithmetic reasoning, code generation, and safety tasks demonstrate the highest compression sensitivity, characterized by precipitous performance declines. Figure 5 illustrates the detailed performance impact of various KV cache compression methods across different tasks. This degradation is most pronounced in arithmetic reasoning (c), where performance deteriorates significantly below the compression ratio of 20%, with precision dropping from approximately 0.75 to below 0.5. Among the evaluated methods, ChunkKV [42] and PyramidKV [24] consistently demonstrate superior stability in most tasks, while StreamingLLM [10] exhibits increased sensitivity to aggressive compression. Additionally, R1-Arithmetic reasoning (f) indicates that reasoning LLMs demonstrate enhanced robustness to KV cache compression. Highlighting World Knowledge and Common Sense Reasoning as the most robust tasks, indicating that these tasks are less sensitive to KV cache compression.
@@
Model-Type Robustness: Multi-step reasoning LLMs are more robust to KV cache compression. Figure 8 presents a comparative analysis of LLaMA-3.1-8B across its base (w/o instruct tuned), instruct-tuned, and DeepSeek-R1 distilled variants, illustrating their averaged performance in five compression methods with confidence intervals. Although all three variants exhibit performance degradation at low compression ratios, their degradation trajectories differ significantly. The R1 distilled model demonstrates superior stability, maintaining performance around 0.60 even at a compression ratio 10%. Although the instruct-tuned model achieves a higher initial accuracy (0.8), it exhibits heightened compression sensitivity, with performance deterioration beginning at 30% compression ratio and declining sharply to approximately 0.5 at 10% ratio. These findings suggest that while multi-step reasoning LLMs demonstrate enhanced robustness to KV cache compression, and instruct-tuning improves overall model performance, the latter may inadvertently increase model vulnerability to aggressive compression, particularly at compression ratios below 30%.
(a) Many-shot Arithmetic Reasoning on LLaMA3.1-8B-Instruct
(b) Many-shot Arithmetic Reasoning on DeepSeek-R1-Distill-Llama-8B
Figure 6: Many-shot scenario on KV cache compression
Observation 3.
Prompt Length Vulnerability: Shorter prompts are more vulnerable to KV cache compression. As illustrated in Figure 8, the effect of KV cache compression is markedly different with varying prompt lengths (shot numbers). Scenarios with fewer shots (for example, one-shot and two-shot) demonstrate heightened sensitivity to compression; their performance degrades more precipitously below a compression ratio of 30% compared to scenarios with a greater number of shots (e.g., 4-8 shots). For example, in 1-shot settings, performance decreases from 0.5 to 0.05 as the compression ratio decreases from 30% to 10%. In contrast, 8-shot settings experience a less severe reduction, from 0.75 to 0.5, under the same compression conditions. This suggests that prompts with more shots, by virtue of containing more contextual examples, offer a richer set of reference points for the model. Consequently, the model’s reliance on any single example being perfectly preserved in the compressed KV cache is reduced, leading to greater robustness against aggressive compression.
Observation 4.
Chunk-Level Superiority: Chunk-level compression is more effective for long-context structured reasoning tasks. Inspired by Agarwal et al. [44], we consider many-shot in-context learning as a long-context reasoning task, which is more complex than existing long-context benchmarks, such as LongBench and NIAH. Figure 6 shows the performance of KV cache compression methods on a 50-shot GSM8K task, where the prompt length exceeds 4K tokens. From the figure, we observe that ChunkKV [42] demonstrates the most stability when the compression ratio is below 10% on both LLaMA-3.1-8B-Instruct and DeepSeek-R1-Distill-Llama-8B, indicating that in more complex long-context arithmetic reasoning tasks, chunk-level retention is more effective at preserving semantic information. Highlighting the effectiveness of chunk-level compression for long-context structured reasoning tasks.
Figure 7: Performance Comparison of KV Cache Compression Methods on different training dynamics on Arithmetic Reasoning
@@
Prompt-Gain Sensitivity: KV cache compression significantly reduces performance gains from ICL and CoT. As shown in Table 2, different tasks exhibit varying levels of performance improvement from zero-shot to CoT prompting. Arithmetic reasoning shows a dramatic improvement of 50.41%, while World Knowledge demonstrates a more modest gain of 6.20%.
@@
Arithmetic Reasoning $29.04$ $79.45$ $+50.41$
World Knowledge $62.62$ $68.82$ $+6.20$
From Figure 4, we find that tasks with larger CoT improvements, such as Arithmetic reasoning, are more sensitive to KV cache compression. This suggests that when a task is heavily based on CoT to achieve better performance, compression of these crucial prompt elements has a more substantial impact on model performance. In contrast, tasks like World Knowledge, where the performance gain from CoT is smaller, show more resilience to KV cache compression, likely because the model relies more on its inherent knowledge than on the specific examples in the prompt.
Table 3: KV cache compression methods’ performance on LG-GSM8K
@@
Observation 6.
Long-Context Generation Sensitivity: KV cache compression exhibits significant performance degradation in long-context generation tasks. As demonstrated in Table 3, our evaluation of three unified compression methods—StreamingLLM, H2O, and PyramidInfer—on LG-GSM8K reveals substantial performance limitations. In this arithmetic reasoning task with approximately 4k token generation duration, compression methods show notable deterioration, with performance declining by more than 20% at compression ratios below 30%. The ShotKV is our proposed method that aims to improve the performance of KV cache compression on Long-Context Generation tasks, details in Section 4.